# Muon beams towards muonium physics: progress and prospects

### v1

@@ V.2 Application of negative muon
During implantation, the loses most of its energy through electromagnetic interactions with the target material. Once sufficiently slowed, it is captured by the nucleus, forming a muonic atom and immediately entering a rapid cascade down to the 1S ground state as shown in Fig. 30. Using the Bohr model [PhysRev.177.164, PhysRevC.71.058501], this cascade can be described in terms of atomic radii, where the radius of the orbit with quantum number is given by The radius of the orbit with the quantum number can be expressed as
@@
where the energy is about 207 times larger than the corresponding electron energy . During the cascade process, the rapidly transitions to the lower atomic energy levels; the primary energy emission is in the form of characteristic X-rays or, in some cases, Auger electrons. The energies of these muonic X-rays typically range from several up to a few , and are highly element-specific due to their dependence on the atomic number . This property makes muonic X-ray emission a powerful tool for elemental analysis and depth profiling in material science. Analogous to the empirical Moseley’s law [moseley1913xciii], which describes the relationship between atomic number and X-ray emission energies for electronic transitions, the energies of muonic X-ray transitions can be similarly expressed as
@@ V.2.2 Muon Induced X-ray Emission (MIXE) technique
Muonic Induced X-ray Emission (MIXE) is a nondestructive technique for elemental analysis, particularly in materials science. By utilizing high-energy X-rays emitted during the muon-nucleus interactions, MIXE enables depth-dependent characterization of elemental composition, These features make MIXE potentially suitable for specialized applications in fields such as biology and archaeology, where nondestructive analysis is crucial.
@@
A schematic diagram of the PSI MIXE spectrometer is shown in Fig. 31, illustrating a representative configuration of the spectrometer frame, detection unit, and sample environment. The surrounding HPGe array ensures a high energy resolution for muon-induced X-rays. The characteristic X-ray energies enable element identification, while their intensities provide quantitative information on elemental composition.
@@ V.3 Perspectives on muon applications
Recent progress on muon source and detector technology has broadened the application of muon science from a limited field to a research platform covering varying research areas. Positive muon have become an essential tool for the study of local magnetism, superconductivity, and chemical reaction mechanisms, represented by the SR and muonium chemistry. Since the muon spin procession can reflect the weak internal magnetic field and dynamic fluctuation, it fills in the gap left by other spectroscopic techniques such as NMR and neutron scattering. On the other hand, negative muon techniques, such as MIXE and SR, offer complementary capabilities, enabling non-destructive elemental analysis and the investigation of hydrogen diffusion in complex environments.
With the construction of next-generation muon facilities, the beam intensity and quality will be further improved. These superior beam performances will bring further challenges and opportunities for the muon applications. This evolution will not only refine our understanding of condensed matter systems but also open up novel avenues in archaeology, battery research, and fundamental physics. Meanwhile, the high-intensity muon beams will bring new challenges for the detector technology. For example, the high counting rate may lead to the pile-up effect, which will significantly deteriorate the detector performance. Therefore, the development of new detector technology is essential to fully exploit the advantages of high-intensity muon beams. In conclusion, the combination of advanced beams and next-generation detectors promises a bright future for muon applications.
@@ VI Summary and outlook
Research on various scenarios involving accelerator muon beams has continued for nearly half a century. The demand for high-intensity muon beams is rapidly increasing. Therefore, several countries, especially China, are currently planning and constructing next-generation muon facilities that can drive future experiments to higher precision. In summary, muonium, as a unique system, can support physics motivations across particle physics, nuclear physics, and materials science. Understanding of fundamental symmetries in nature also drives innovation in beam technology, particle detection, and advanced computing. By leveraging world‑class infrastructure, muon beams towards muonium physics across the road exemplifies how fundamental research can stimulate technological progress and international collaboration. With the construction of new muon facilities, the studies of muonium are expected to constantly provide exciting contributions to the particle physics community and the related areas.

### v2

@@ V.2 Application of negative muon
During implantation, the loses most of its energy through electromagnetic interactions with the target material. Once sufficiently slowed, it is captured by the nucleus, forming a muonic atom and immediately entering a rapid cascade down to the 1S ground state as shown in Fig.˜30. Using the Bohr model [PhysRev.177.164, PhysRevC.71.058501], this cascade can be described in terms of atomic radii, where the radius of the orbit with quantum number is given by
@@
where the energy is about 207 times larger than the corresponding electron energy . During the cascade process, the rapidly transitions to the lower atomic energy levels; the primary energy emission is in the form of characteristic X-rays or, in some cases, Auger electrons. The energies of these muonic X-rays typically range from several up to a few , and are highly element-specific due to their dependence on the atomic number . This property makes muonic X-ray emission a powerful tool for elemental analysis and depth profiling in materials science. Analogous to the empirical Moseley’s law [moseley1913xciii], which describes the relationship between atomic number and X-ray emission energies for electronic transitions, the energies of muonic X-ray transitions can be similarly expressed as
@@ V.2.2 Muon Induced X-ray Emission (MIXE) technique
MIXE is a nondestructive technique for elemental analysis, particularly in materials science. By utilizing high-energy X-rays emitted during the muon-nucleus interactions, MIXE enables depth-dependent characterization of elemental composition. These features make MIXE potentially suitable for specialized applications in fields such as biology and archaeology, where nondestructive analysis is crucial.
@@
A schematic diagram of the PSI MIXE spectrometer is shown in Fig.˜31, illustrating a representative configuration of the spectrometer frame, detection unit, and sample environment. The surrounding HPGe array ensures a high energy resolution for muon-induced X-rays. The characteristic X-ray energies enable element identification, while their intensities provide quantitative information on elemental composition.
@@ V.3 Perspectives on muon applications
Recent progress on muon source and detector technology has broadened the application of muon science from a limited field to a research platform covering varying research areas. Positive muons have become an essential tool for the study of local magnetism, superconductivity, and chemical reaction mechanisms, represented by the SR and muonium chemistry. Since the muon spin precession can reflect the weak internal magnetic field and dynamic fluctuation, it fills in the gap left by other spectroscopic techniques such as NMR and neutron scattering. On the other hand, negative muon techniques, such as MIXE and SR, offer complementary capabilities, enabling non-destructive elemental analysis and the investigation of hydrogen diffusion in complex environments.
With the construction of next-generation muon facilities, the beam intensity and quality will be further improved. These superior beam performances will bring further challenges and opportunities for the muon applications. This evolution will not only refine our understanding of condensed matter systems but also open up novel avenues in archaeology, battery research, and fundamental physics. Meanwhile, the high-intensity muon beams will bring new challenges for the detector technology. For example, the high counting rate may lead to the pile-up effect, which will significantly degrade the detector performance. Therefore, the development of new detector technology is essential to fully exploit the advantages of high-intensity muon beams. In conclusion, the combination of advanced beams and next-generation detectors promises a bright future for muon applications.
@@ VI Summary and outlook
Research on various scenarios involving accelerator muon beams has continued for nearly half a century. The demand for high-intensity muon beams is rapidly increasing. Therefore, several countries, especially China, are currently planning and constructing next-generation muon facilities that can drive future experiments to higher precision. In summary, muonium, as a unique system, can support physics motivations across particle physics, nuclear physics, and materials science. Understanding of fundamental symmetries in nature also drives innovation in beam technology, particle detection, and advanced computing. By leveraging world-class infrastructure, muon beams towards muonium physics across the road exemplifies how fundamental research can stimulate technological progress and international collaboration. With the construction of new muon facilities, the studies of muonium are expected to constantly provide exciting contributions to the particle physics community and the related areas.